OpenStudy (anonymous):

One Algebra 2 question? I give medals!:)

OpenStudy (anonymous):

Please write out the steps line by line for me, I really need the help! Solve the system of equations. 3x-y=6 6x+y=21

OpenStudy (anonymous):

@uri can you help me with this problem?

OpenStudy (anonymous):

:( anyone...?

OpenStudy (anonymous):

Darn... Oh well.. Just turned it in blank. :/

OpenStudy (anonymous):

These are simultaneous equations. You need to take one away from the other or add them together so that one of the unknowns disappears. In your case there a -y in eq1 and a y in eq2, so if you add eq1 and eq2 together the ys will cancel out -y+y=0 3x-y+6x+y=6+21 9x=27 x=27/9 x=3 Then you can just substitute for x in either equations: 3x-y=6 3*3-y=6 3*3-6=y y=3 There :)

OpenStudy (anonymous):

Thank you :)